what are the important postulates of Bohrs atomic structure and rutherford atomic structure

Grade:9

2 Answers

A M S ARUN KRISHNA
8 years ago
BOHR’S MODEL OF ATOM
 
To overcome the above defects of Rutherford’s model, Niels Bohr in 1913 gave a modification based on Quantum theory of radiation. The important postulates are:
  • The electrons revolve round the nucleus only in certain fixed energy levels  called orbits. These orbits are associated with definite energies and are called energy shells or energy levels or quantum levels. These are numbered as 1, 2, 3, 4 ….. etc. (starting from the nucleus) are designated as K, L, M, N ….etc. (Fig. 3.2).
  • As long as an electron remains in a particular orbit, it does not lose or gain energy. This means that energy of an electron in a particular path remains constant. Therefore, these orbits are also called stationary states.
  • If an electron jumps from one stationary state to another, it will absorb or emit radiation of a definite frequency giving a spectral line of that frequency which depends upon the initial and final levels. When an electron jumps back to the lower energy level, it radiates same amount of energy in the form of radiation.
`The exchange of energy is possible only when electron jumps from one energy level to another energy level.
  E = E2 – E1 = hv = hc/ʎ  joule
Where h is planck’s constant having fixed value
                h = 6.63 × 10-34 J / second  and ʎ is wavelength.
                When an electron falls from an orbit of high energy level to lower energy level, the difference in energy is radiated in the form of electromagnetic radiation of particular wavelength. Since each atom has its specific energy levels, it can energy levels, it can emit radiations of specific wavelength. Ground state means the lowest energy state.  When the electrons absorb energy and jump to outer orbits, this state is called excited state.
HOW ARE ELECTRONS DISTRIBUTED IN DIFFERENT SHELLS
 
              The distribution of electrons into different orbits of an atom was suggested by Bohr and Bury. The following rules are followed for writing the number of electrons in different energy levels or shells:
  • The maximum number of electrons present in a shell is given by the formula 2n2, where ‘n’ is the orbit number or energy level index, 1,2,3,…. Hence the maximum number of electrons in different shells are as follows:  
first orbit or K-shell will be = 2 × 12 = 2, second orbit or L-shell will be = 2 × 22 = 8, third orbit or M-shell will be = 2 × 32 = 18, fourth orbit or N-shell will be  2 × 42= 32, and so on.
  • The maximum number of electrons that can be accommodated in the outermost orbit is 8.
  • Electrons are not accommodated in a given shell, unless the inner shells are filled. That is, the shells are filled in a step-wise manner.
 
LIMITATIONS OF BOHR’S THEORY
 
  • According to Bohr, the radiation results when an electron jumps from one energy orbit to another energy orbit, but how this radiation occurs is not explained by Bohr.
  • This theory was applicable only for monoelectronic system that is H, He+, Li++ and H2+
  • Bohr Theory had explained the existence of various lines in H spectrum, but it predicted that only a series of lines exist. At that time this was exactly what had been observed. However, as better instruments and techniques were developed, it was realized that the spectral line that had been thought to be a single line was actually a collection of several lines very close together (known as fine spectrum). Thus for example, the single H¥-spectral line of Balmer series consists of many lines very close to each other.
  • It fails to explain why spectrum of hydrogen atom is discontinuous.
  • Thus the appearance of the several lines implies that there are several energy levels, which are close together for each quantum number n. This would require the existence of new quantum numbers.
  • Bohr’s theory has successfully explained the observed spectra for hydrogen atom and hydrogen like ions (e.g. He+, Li2+, Be3+ etc.), it cannot explain the spectral series for the atoms having a large number of electrons.
  • Bohr assumes that an electron in an atom is located at a definite distance from the nucleus and is revolving round it with definite velocity, i.e. it is associated with a fixed value of momentum. This is against the Heisenberg’s Uncertainty Principle according to which it is impossible to determine simultaneously with certainty the position and the momentum of a particle.
  • No explanation for Zeeman effect: If a substance which gives a line emission spectrum, is placed in a magnetic field, the lines of the spectrum get split up into a number of closely spaced lines. This phenomenon is known as Zeeman effect. Bohr’s theory has no explanation for this effect.
  • No explanation of the Stark effect: If a substance which gives a line emission spectrum is placed in an external electric field, its lines get spilt into a number of closely spaced lines. This phenomenon is known as Stark effect. Bohr’s theory is not able to explain this observation as well. 
please approve my answer
deapan
8 years ago
RUTHERFORD’S MODEL OF AN ATOM
 
Rutherford designed an experiment to find the model of the atom called α - particle scattering experiment in 1911.
 
In this experiment, a stream of high energy alpha particles emanating from radium was directed at a thin foil of gold metal. The thin gold metal foil had a circular fluorescent zinc sulphide screen around it. Whenever alpha particles struck the screen, a tiny flash of light was produced at that point.
The gold foil was about 1000 atoms thick approximately thickness of 100 nanometre. Alpha particles are doubly-charged helium ions and highly energetic particles. Since they have a mass of 4 u, the fast-moving α-particles have a considerable amount of energy. It was expected that alpha particles would be deflected by the sub-atomic particles in the gold atoms. But, the particle scattering experiment gave totally unexpected results. The following observations were made:
  • Most of the fast moving particles passed straight through the gold foil.
  • Some of the particles were deflected by the foil by small angles.
  • One out of every 10000 particles reflected about 1800.
Rutherford concluded from the a particle scattering experiment that
 
The atom is mostly composed of empty space. Very few alpha particles were deflected by 1800. The deflection must be due to enormous repulsive force showing that positive charge of the atom is not spread throughout the atom. The entire positive charge and most mass of the atom has to be densely concentrated in a very small volume known as nucleus that repelled and deflected the positively charged alpha particles.
The diameter of the nucleus is estimated by Rutherford as 10-15 cm in contrast to that of an atom to be 10-10 cm. The nucleus is surrounded by electrons that move around the nucleus with very high speed in circular paths called orbits. Electrons and nucleus are held together by the electrostatic forces of attraction
please approve my answer